# Performant Automatic BLAS Offloading on Unified Memory Architecture with OpenMP First-Touch Style Data Movement

### v1

@@ 1. Introduction
In our previous work(Li et al., 2024; Li and Wang, 2024), we presented a proof-of-concept framework for doing symbol interception and replacement with Dynamic Binary Instrumentation (DBI), and preliminary test of automatically offload the *gemm (general matrix multiplication) routines, proofing that performat automatic BLAS offload is achievable. In this paper, I further extend the implementation to all level-3 BLAS calls, and focus on discussing a novel data management strategy named Device First-use policy. Inspired by the OpenMP First-Touch memory management approach in multi-socket CPU or NUMA programming, The Device First-Use policy is designed for CPU-GPU systems where data is moved to GPU memory (or general device memory) upon its first use by a GPU kernel. This approach is both simple and effective, minimizing data transfers in practical BLAS use cases. Tests across several BLAS-intensive scientific applications on up to hundreds of GPU nodes demonstrate significant speedups. Although all tests are done on the NVIDIA Grace-Hopper system, the methodology is generic for any CPU-GPU system with cache coherency.
The remainder of this paper is organized as follows: Section 2 reviews the NVIDIA Grace-Hopper unified memory architectures and related BLAS offloading work. Section 3 details the implementation of SCILIB-Accel. In Section 4, we apply the tool to two BLAS-intensive scientific computing codes on up to 200 Grace-Hopper nodes and discuss the results along with performance issues of the NVIDIA Grace-Hopper system. Finally, Section 5 concludes the paper.
@@ 2.2. Previous automatic BLAS offload attempts
Numerous attempts have been made to automatically accelerate CPU BLAS calls since the early adoption of GPUs in HPC. Cray LIBSCI_ACC(Hewlett Packard Enterprise, [n. d.]; Poxon, 2013), available for over a decade, was deployed on the Titan supercomputer for NVIDIA Tesla K20 GPU, supporting selected BLAS, LAPACK, and ScaLAPACK routines for offload when the library module is loaded. Similarly, IBM ESSL(IBM, [n. d.]) is capable of automatically offload selected BLAS, LAPACK, and FFTW calls but requires the accelerated version of math library, libesslsmpcuda, is linked. NVIDIA’s NVBLAS(NVIDIA, 2024) serves as a drop-in replacement for CPU BLAS calls, allowing users to configure host BLAS libraries and selected routines for offload. By $LD\_PRELOAD$ NVBLAS, dynamically linked CPU BLAS is replaced without relinking. Unfortunately, the NVBLAS tool is heavily over-engineered, it uses the cuBLASXT as the backend instead of cuBLAS and has an acceptable overhead(Li et al., 2024).
These tools make offload decisions at runtime based on workload sizes and handles data movement automatically. Overall, these libraries are tailored for conventional GPU architectures, and frequent data movement is unavoidable, therefore suffers poor performance for small and medium sized matrix math in real workloads.
@@ 3. Performant Automatic BLAS Offload
Figure 1. Workflow of Automatic BLAS Offload
A basic workflow of automatic BLAS offload tool is illustrated in fig:workflow where an dgemm call in the caller code is intercepted and redirected to a BLAS wrapper that manages data movement and makes the GPU BLAS call. Conceptually this workflow contains two tasks: 1) intercept BLAS symbols and replace it with a BLAS wrapper where GPU BLAS is called, 2) manage data movement between CPU and GPU resident memories. In the following content, we will discuss how these tasks are implemented in SCILIB-Accel, most critially how the data movement can be optimized using a novel data movement strategy inspired by the OpenMP First-Touch data placement policy.
@@ 3.1. Symbol Interception
For automatic BLAS offload, we intercept BLAS calls where the trampoline function (BLAS wrapper function) maintains the same signature as the original function. This results in minimal overhead, as the register data for function arguments remain undisturbed. This mechanism finds extensive use in profilers, and here we use the PEAK(Wang and Li, 2023) lightweight profiler we have developed as a DBI framework, ensuring portability across various architectures including but not limited to x86 and ARM. Inside the DBI framework, the FRIDA-GUM(Frida, nd) binary instrumentation backbend is chosen for simplicity and performance.
The SCILIB-Accel automatic offload library can be attached to user application by $LD\_PRELOAD$ the library. The SCILIB-Accel initialization function is placed into the .init_array section of Linux ELF to search and replace BLAS symbols along with other initialization tasks such as setting GPU memory pool, initialize cuBLAS, etc. Similarly, the SCILIB-Accel finalization function is inserted into the .fini_array section of the ELF to collect statistics and handle clean up tasks.
@@
As DBI is widely used in profilers, the DBI symbol interception in SCILIB-Accel can cause conflicts when doing profiling. To be profiler friendly, an implementation of SCILIB-Accel using dlsym() to dynamically resolve shared library symbols to intercept BLAS calls is also provided. This approach works by defining the wrapper function to be the same name as the function to be intercepted, and by prepending the SCILIB-Accel library in $LD\_PRELOAD$, the symbols in the wrappers get used and the original function symbol can be obtained by looking up the next available symbol using dlsym(). This dlsym-based version have no issue being used with profilers, but can only intercept dynamically linked BLAS.
@@ 3.2. Data Movement Strategies
Managing data movement is often the most critical part of GPU porting as data transfer speed has been a limiting factor. It is even more so for developing automatic offload tool as the tool deals with pure CPU code that is totally untuned for GPU. Here, we discuss three data movement strategies that utilize different features of Grace-Hopper and it helps us better understand the challenge and opportunities for doing automatic offload on unified memory architecture.
@@ 3.2.1. Strategy 1, Mem-Copy
This is the most intuitive strategy and used by other tools. The pseudocode is listed below. Upon interception of a BLAS call and code is redirected to a BLAS wrapper, the input matrices are copied from host memory to GPU memory, then resultant matrix is copied back after cuBLAS execution. then resultant matrix is copied back after cuBLAS execution. This strategy works on all GPUs including the conventional PCIe-based cards without needing unified memory capability, This strategy works on all GPUs including the conventional PCIe-based cards without needing unified memory capability, but at the cost of frequent data movement. This strategy can be effective for codes handling very large matrices where the compute time far exceeds the data transfer cost, but won’t be useful for most codes that only runs small to medium sized matrices in practical use. This policy is studied here mostly for helping us understand the limitation of the conventional automatic BLAS offload approach in all other existing tools.
@@ 3.2.2. Strategy 2, counter-based migration
Since CPU resident memory (LPDDR5X) and the GPU resident memory (HBM3) are physically unified with cache-coherent NVLink C2C, CPU matrix pointers can be passed directly to cuBLAS calls. One can also use numactl to force all memory resident on the HBM, but given the poor bandwidth of CPU access HBM, a performance penalty is expect. A new feature in Grace-Hopper designed to better serve the unified memory is the access counter on Hopper GPU, the CUDA runtime will automatically move memory page from LPDDR5X to HBM3 based on remote memory access detected by the counter. This counter-based migration mechanism can serve automatic offload when we pass CPU resident matrices to GPU kernel. A pseudocode example is shown below.
@@ 3.2.3. Strategy 3, Device First-Use policy
Since the BLAS wrapper used to replace the CPU BLAS calls only knows the memory addresses of the matrices, implementing Device First-Use policy requires move data from the CPU resident memory to device (GPU) resident memory without reallocation or disrupting to the virtual memory address used by the CPU binary. This can be achieved by relocating the physical memory page and updating the page table to remap the virtual memory to new physical memory locations. Figure 2 illustrates the working of virtual memory in modern operation system and how physical memory page can be dynamically reassigned.
@@
Figure 2. Virtual Memory and Physical Page Migration
@@
How does this policy improve data reusability for the GPU? To answer this, we need to examine how BLAS is typically used in scientific applications. In most cases, scientific problems are not solved with a single BLAS call. Instead, they involve a sequence of BLAS operations, such as $C=A\times B$ followed by $E=D\times C$, and so on. In these workflows, intermediate matrices (e.g., C) are frequently reused in subsequent operations. Data reuse is also particularly common in block matrix multiplications, where each block of a matrix is multiplied by multiple blocks from another matrix. Additionally, many scientific codes adopts an iterative approach, such as the self-consistent field process to solve partial differential equations in quantum chemistry or physics, the same matrices and memory pointers are re-used across all the iterations. All the data only need to be moved to the GPU once and can be re-used by subsequent iterations. All these common use cases conceptually justifies the appropriateness of Device First-Use policy.
@@ 3.3. Usage Instructions
The automatic BLAS offload wrappers are compiled and linked together as a shared library file (.so). All users need to do is to load the SCILIB-Accel library by $LD\_PRELOAD$ the .so file as shown below, then run their CPU binary as normal.
@@
Several optional environmental variables can be set to tweak offload behaviors including:
@@ 4. Performance Testings and Discussion
In this section, I perform application tests using production HPC codes. Since quantum chemistry/physics are known to heavily relying on BLAS operations as linear algebra is the natural language of quantum mechanics, two codes, MuST and PARSEC, from this scientific domain are chosen, and they also part of the Characteristic Science Application (CSA)(Foundation, 2021) efforts funded the National Scientific Foundation (NSF) as representative workloads for the Leadership Class Computing Facility.
Different offload strategies are extensively tested to showcase the performance of Device First-Use data movement policy, and understand limitations of data movement approaches. In all test cases, the offload threshold is $N_{avg}>500$ (see Section 3.3 for details) will be offloaded which is proven to be appropriate for these applications on Grace-Hopper.
It is important to emphasize that all performance comparisons presented here are based on an equal number of nodes: Grace-Grace nodes (two CPUs with 144 cores) and Grace-Hopper nodes (one CPU and one GPU). This approach ensures both simplicity and fairness in the analysis. Power and cost are also key factors in performance study. The power consumption of a Grace-Hopper node is approximately twice that of a Grace-Grace node under full load. Additionally, for each node-hour, the Service Unit (SU) charged to users for Grace-Hopper nodes is three times higher than that for Grace-Grace nodes. This charge rate reflects the costs associated with acquiring, maintaining, and supporting these different types of nodes.
@@ 4.1. Test Environment
  • Grace-Hopper Nodes: Each node is equipped with one 72-core Grace CPU (120 GB LPDDR5X memory) and one H100 GPU (96 GB HBM3 memory). The Grace-Hopper superchip is power-capped at 900W.
  • Grace-Grace Nodes: Each node features two 72-core Grace CPUs with a combined Thermal Design Power (TDP) of 500W.
@@ 4.2. Application Test 1: MuST
Figure 3. MuST: Strong Scaling Test for CPU Run and Automatic GPU Offload Run
MuST (Multiple Scattering Theory)(Wang et al., 1995; Eisenbach et al., 2017) is a package designed to perform electronic structure calculations, it solves the Kohn-Sham equation by solving the Green’s function. The LSMS calculation method is designed for large systems with linear scalability to the system size. This method is won the 2009 Gordon Bell prize (Eisenbach et al., 2009). The code has a heavy dependency on BLAS operations, mostly zgemm and ztrsm, which often exceeds 80% of runtime on CPU. A major portion of the BLAS calls are from LAPACK routines zgetrs and zgetrf. Most matrices are squared or near-square shaped. MuST natively supports GPU offload, the native method implemented in MuST offloads the matrix inverse onto a GPU by calling cuSOLVER.
The test workload calculates energy of a CoCrFeMnNi supercell alloy using the LSMS method. Total atom number in the supercell is 5600 and concentration of each element is identical. The number of energy grid is 32. The calculation is limited to 3 self-consistent steps to reduce benchmark cost.
@@
MuST is thoroughly tested at large scale on CPU, on GPU through automatic offload and native CUDA port. Table 3 summarizes the performance comparison of different run strategies on 50 Grace-Grace CPU nodes or 50 Grace-Hopper nodes. As mentioned before, the code strongly relying BLAS operations, the two major BLAS routines zgemm and ztrsm consumes about 2080s out of the total 2318s runtime. Using the native CUDA port from the developers, about 1.4x speedup is achieved comparing to CPU run. Surprisingly, all auto offload strategies are faster than the native CUDA code illustrating the complexity and challenge of CUDA programming, the developers will need spend substantial more amount porting efforts to polish their CUDA code. With the most basic data management method that copies (cudaMemcpy) matrices to/from GPU for every cuBLAS call, the total runtime reduces to 1098s, but 292s are spent in just moving the data around. This reaffirms that optimizing data movement is still critical on Grace-Hopper even with the fast NVLink-C2C interconnect. The counter-based migration approach works okay, total runtime is better than doing frequent cudaMemcpy. Note that the page migration time is included in the BLAS call time as the counter-based does the data movement automatically behind the scene with GPU kernel. The novel Device First-Use policy is substantially better other approaches, total runtime is reduced to 824s, about 2.8x faster than the CPU run, the total data movement time is reduced to just 4.8s. More analysis shows that under Device First-Use policy, each matrix that gets migrated to the GPU resident memory gets reused 780 times by subsequent BLAS calls. Such high level of data reusage is the key factor of the good performance we achieve here. Also note that the BLAS (zgemm and ztrsm) time in Device First-Use is much longer than the corresponding time in Mem-Copy policy, this is due to a performance issue of CUDA kernel accessing GPU memory allocated by system malloc, more details are discussed in Section 4.4.3. If such performance issue is resolved by NVIDIA, performance of automatic offload can be further improved.
Large scale strong scaling tests were also performed for this code. Table 4 shows the runtime of CPU run, GPU run with native CUDA code, and automatic offload using Device First-Use policy. Scaling range goes from 25 nodes to 200 nodes. Throughout the tests, the automatic offload approach is consistent 2x faster than the native CUDA code, and up to 3x faster than the CPU run, break even with the extra node-hour charge rate for GPU so users running on GPU not only gets faster time-to-solution but also more efficient in terms of energy consumption and cost. The strong scaling data is visualized in Figure 3, both the CPU run and automatic offload GPU run have excellent scability, very close to linear scaling at this wide test range.

### v2

@@ 1. Introduction
In my previous work(Li et al., 2024; Li and Wang, 2024), I presented a proof-of-concept framework for doing symbol interception and replacement with Dynamic Binary Instrumentation (DBI), and preliminary tests of automatically offloading the *gemm (general matrix multiplication) routines, proving that performat automatic BLAS offload is achievable. In this paper, I further extend the implementation to all level-3 BLAS calls, and focus on discussing a novel data management strategy named Device First-use policy. Inspired by the OpenMP First-Touch memory management approach in multi-socket CPU or NUMA programming, the Device First-Use policy is designed for CPU-GPU systems where data is moved to GPU memory (or general device memory) upon its first use by a GPU kernel. This approach is both simple and effective, minimizing data transfers in practical BLAS use cases. Tests across several BLAS-intensive scientific applications on up to hundreds of GPU nodes demonstrate significant speedups. Although all tests are done on the NVIDIA Grace-Hopper system, the methodology is generic for any CPU-GPU system with cache coherency.
The remainder of this paper is organized as follows: Section 2 reviews the NVIDIA Grace-Hopper unified memory architectures and related BLAS offloading work. Section 3 details the implementation of SCILIB-Accel. In Section 4, I apply the tool to two BLAS-intensive scientific computing codes on up to 200 Grace-Hopper nodes and discuss the results along with performance issues of the NVIDIA Grace-Hopper system. Finally, Section 5 concludes the paper.
@@ 2.2. Previous automatic BLAS offload attempts
Numerous attempts have been made to automatically accelerate CPU BLAS calls since the early adoption of GPUs in HPC. Cray LIBSCI_ACC(Hewlett Packard Enterprise, [n. d.]; Poxon, 2013), available for over a decade, was deployed on the Titan supercomputer for NVIDIA Tesla K20 GPU, supporting selected BLAS, LAPACK, and ScaLAPACK routines for offload when the library module is loaded. Similarly, IBM ESSL(IBM, [n. d.]) is capable of automatically offload selected BLAS, LAPACK, and FFTW calls but requires the accelerated version of math library, libesslsmpcuda, linked. NVIDIA’s NVBLAS(NVIDIA, 2024) serves as a drop-in replacement for CPU BLAS calls, allowing users to configure host BLAS libraries and selected routines for offload. By $LD\_PRELOAD$ NVBLAS, dynamically linked CPU BLAS is replaced without relinking. Unfortunately, the NVBLAS tool is heavily over-engineered, it uses the cuBLASXT as the backend instead of cuBLAS and has an acceptable overhead(Li et al., 2024).
These tools make offload decisions at runtime based on workload sizes and they handle data movement automatically. Overall, these libraries are tailored for conventional GPU architectures, where frequent data movement is unavoidable, therefore they suffer poor performance for small and medium sized matrix math in real workloads.
@@ 3. Performant Automatic BLAS Offload
Figure 1. Workflow of Automatic BLAS Offload. The BLAS in the CPU binary is intercepted and replaced with a BLAS wrapper where GPU BLAS call is maded and data is moved between CPU resident memory and GPU resident memory.
A basic workflow of automatic BLAS offload tool is illustrated in Figure 1 where an dgemm call in the caller code is intercepted and redirected to a BLAS wrapper that manages data movement and makes the GPU BLAS call. Conceptually this workflow contains two tasks: 1) intercept BLAS symbols and replace them with a BLAS wrapper where GPU BLAS is called, 2) manage data movement between CPU and GPU resident memories. In the following content, I will discuss how these tasks are implemented in SCILIB-Accel, most critially how the data movement can be optimized using a novel data movement strategy inspired by the OpenMP First-Touch data placement policy.
@@ 3.1. Symbol Interception
For automatic BLAS offload, I intercept BLAS calls where the trampoline function (BLAS wrapper function) maintains the same signature as the original function. This results in minimal overhead, as the register data for function arguments remain undisturbed. This mechanism finds extensive use in profilers, and here I use the PEAK(Wang and Li, 2023) lightweight profiler that was developed by me and my collaborator as a DBI framework, ensuring portability across various architectures including but not limited to x86 and ARM. Inside the DBI framework, the FRIDA-GUM(Frida, nd) binary instrumentation backbend is chosen for simplicity and performance.
The SCILIB-Accel automatic offload library can be attached to the user application by preloading ($LD\_PRELOAD$) the SCILIB-Accel shared library file. The SCILIB-Accel initialization function is placed into the .init_array section of Linux ELF to search and replace BLAS symbols along with other initialization tasks such as setting GPU memory pool, initializing cuBLAS, etc. Similarly, the SCILIB-Accel finalization function is inserted into the .fini_array section of the ELF to collect statistics and handle clean up tasks.
@@
As DBI is widely used in profilers, the DBI symbol interception in SCILIB-Accel can cause conflicts when doing profiling. To be profiler friendly, an implementation of SCILIB-Accel using dlsym() to dynamically resolve shared library symbols to intercept BLAS calls is also provided. This approach works by defining the wrapper function to be the same name as the function to be intercepted, and by prepending the SCILIB-Accel library in $LD\_PRELOAD$, the symbols in the wrappers get used and the original function symbol can be obtained by looking up the next available symbol using dlsym(). This dlsym-based version has no issue being used with profilers, but can only intercept dynamically linked BLAS.
@@ 3.2. Data Movement Strategies
Managing data movement is often the most critical part of GPU porting as data transfer speed has been a limiting factor. It is even more so for developing an automatic offload tool, as the tool deals with pure CPU code that is totally untuned for GPU. Here, I discuss three data movement strategies that utilize different features of Grace-Hopper so that this helps us better understand the challenge and opportunities for doing automatic offload on unified memory architecture.
@@ 3.2.1. Strategy 1, Mem-Copy
This is the most intuitive strategy and used by other tools. The pseudocode is listed below. Upon interception of a BLAS call and redirection to a BLAS wrapper, the input matrices are copied from host memory to GPU memory, and then resultant matrix is copied back after cuBLAS execution. This strategy works on all GPUs including the conventional PCIe-based cards without needing unified memory capability, but at the cost of frequent data movement. Such strategy can be effective for codes handling very large matrices where the compute time far exceeds the data transfer cost, but won’t be useful for most codes that only runs small to medium sized matrices in practical use. This policy is studied here mostly for helping us understand the limitation of the conventional automatic BLAS offload approach in all other existing tools.
@@ 3.2.2. Strategy 2, counter-based migration
Since the CPU resident memory (LPDDR5X) and the GPU resident memory (HBM3) are physically unified with cache-coherent NVLink C2C, CPU matrix pointers can be passed directly to cuBLAS calls. One can also use numactl to force all memory to be resident on the HBM, but given the poor bandwidth of CPU access HBM, a performance penalty is expect. A new feature in Grace-Hopper designed to better serve the unified memory is the access counter on Hopper GPU, namely the CUDA runtime will automatically move memory pages from LPDDR5X to HBM3 based on remote memory accesses detected by the counter. This counter-based migration mechanism can serve automatic offload when CPU resident matrices are passed to a GPU kernel. A pseudocode example is shown below.
@@ 3.2.3. Strategy 3, Device First-Use policy
Since the BLAS wrapper used to replace the CPU BLAS calls only knows the memory addresses of the matrices, implementing Device First-Use policy requires moving data from the CPU resident memory to device (GPU) resident memory without reallocation or disrupting to the virtual memory address used by the CPU binary. This can be achieved by relocating the physical memory page and updating the page table to remap the virtual memory to new physical memory locations. Figure 2 illustrates the working of virtual memory in modern operation systems and how a physical memory page can be dynamically reassigned.
@@
Figure 2. Virtual Memory and Physical Page Migration. Diagram showing translation between virtual and physical memory. The physical memory page could be moved without impacting virtual memory page or the viewpoint of memory from the executing binary side.
@@
How does this policy improve data reusability for the GPU? To answer this, let’s examine how BLAS is typically used in scientific applications. In most cases, scientific problems are not solved with a single BLAS call. Instead, they involve a sequence of BLAS operations, such as $C=A\times B$ followed by $E=D\times C$, and so on. In these workflows, intermediate matrices (e.g., C) are frequently reused in subsequent operations. Data reuse is also particularly common in block matrix multiplications, where each block of a matrix is multiplied by multiple blocks from another matrix. Additionally, many scientific codes adopts an iterative approach, such as the self-consistent field process to solve partial differential equations in quantum chemistry or physics, where the same matrices and memory pointers are re-used across all the iterations. All the data only need to be moved to the GPU once and can be re-used by subsequent iterations. All these common use cases conceptually justify the appropriateness of the Device First-Use policy.
@@ 3.3. Usage Instructions
The automatic BLAS offload wrappers are compiled and linked together as a shared library file (.so). All users need to do is to load the SCILIB-Accel library by $LD\_PRELOAD$ the .so file as shown below, and then run their CPU binary as normal.
@@
Several optional environmental variables can be set to tweak offload behaviors, including:
@@ 4. Performance Testings and Discussion
In this section, I perform application tests using production HPC codes. Since quantum chemistry/physics are known to heavily rely on BLAS operations, as linear algebra is the natural language of quantum mechanics, two codes, MuST and PARSEC, from this scientific domain are chosen. These codes are part of the Characteristic Science Application (CSA)(Foundation, 2021) efforts funded the National Scientific Foundation (NSF) as representative workloads for the Leadership Class Computing Facility.
Different offload strategies are extensively tested to showcase the performance of Device First-Use data movement policy, and understand limitations of data movement approaches. In all test cases, the offload threshold is $N_{avg}>500$ (see Section 3.3 for details), which is proven to be appropriate for these applications on Grace-Hopper.
It is important to emphasize that all performance comparisons presented here are based on an equal number of nodes: Grace-Grace nodes (two CPUs with 144 cores) and Grace-Hopper nodes (one CPU and one GPU). This approach ensures both simplicity and fairness in the analysis. Power and cost are also key factors in performance studies. The power consumption of a Grace-Hopper node is approximately twice that of a Grace-Grace node under full load. Additionally, for each node-hour, the Service Unit (SU) charged to users for Grace-Hopper nodes is three times higher than that for Grace-Grace nodes. This charge rate reflects the costs associated with acquiring, maintaining, and supporting these different types of nodes.
@@ 4.1. Test Environment
  • Grace-Hopper Nodes: Each node is equipped with one 72-core Grace CPU (120 GB LPDDR5X on-board memory) and one H100 GPU (96 GB HBM3 memory). The Grace-Hopper superchip is power-capped at 900W.
  • Grace-Grace Nodes: Each node features two 72-core Grace CPUs with a combined Thermal Design Power (TDP) of 500W. Each Grace chip has 120GB LPDDR5X memory integrated as well.
@@ 4.2. Application Test 1: MuST
Figure 3. MuST: Strong Scaling Test for CPU Run and Automatic GPU Offload Run. Energy calculation of the 5600-atom CoCrFeMnNi alloy system is tested on on various number of Grace-Grace nodes (2 Grace CPUs per node) and Grace-Hopper nodes (1 CPU and 1 GPU per node). Speedup of automatic BLAS offload on GPU nodes is up to 3x comparing to the same number of CPU nodes, breaking even with the cost ratio of node-hour charged for the Vista(Texas Advanced Computing Center, 2024) system at TACC.
MuST (Multiple Scattering Theory)(Wang et al., 1995; Eisenbach et al., 2017) is a package designed to perform electronic structure calculations, it solves the Kohn-Sham equation by solving the Green’s function. The LSMS calculation method is designed for large systems with linear scalability to the system size. This method is won the 2009 Gordon Bell prize (Eisenbach et al., 2009). The code has a heavy dependency on BLAS operations, mostly zgemm and ztrsm, which often exceed 80% of the runtime on CPU. A major portion of the BLAS calls are from the LAPACK routines zgetrs and zgetrf. Most matrices are squared or near-square shaped. MuST natively supports GPU offload through a CUDA implementation and offloads the matrix inverse onto a GPU by calling cuSOLVER.
The test workload calculates the energy of a CoCrFeMnNi supercell alloy using the LSMS method. Total number of atoms in the supercell is 5600 and the concentration of each element is identical. The number of energy grid is 32. The calculation is limited to 3 self-consistent field (SCF) steps in order to reduce benchmark cost.
@@
MuST is thoroughly tested at large scale on CPU, and on GPU through both the automatic offload and the native CUDA port. Table 3 summarizes the performance comparison of different run strategies on 50 Grace-Grace CPU nodes or 50 Grace-Hopper nodes. As mentioned before, the code strongly relies on BLAS operations. In this particular test workload, the two major BLAS routines zgemm and ztrsm consumes about 2080s out of the total 2318s runtime. Using the native CUDA port from the developers, about 1.4x speedup is achieved comparing to the CPU-only execution. Surprisingly, all auto offload strategies are faster than the native CUDA code, illustrating the complexity and challenge of CUDA programming, the developers will need spend substantial more amount porting efforts to polish their CUDA code. With the most basic data management method that copies (cudaMemcpy) matrices to/from GPU for every cuBLAS call, the total runtime is reduced to 1098s, but 292s are spent in just moving the data around. This reaffirms that optimizing data movement is still critical on Grace-Hopper even with the fast NVLink-C2C interconnect. The counter-based migration approach works okay, as the total runtime is better than doing frequent cudaMemcpy. Note that the page migration time is included in the BLAS call time as the counter-based does the data movement automatically behind the scene with GPU kernel. The novel Device First-Use policy is substantially better than other approaches, since the total runtime is reduced to 824s, about 2.8x faster than the CPU run, and the total data movement time is reduced to just 4.8s. More analysis shows that under Device First-Use policy, each matrix that gets migrated to the GPU resident memory gets reused 780 times by subsequent BLAS calls. Such high level of data reusage is the key factor of the good performance that is achieved here. Also note that the BLAS (zgemm and ztrsm) time in Device First-Use is much longer than the corresponding time in Mem-Copy policy. This is due to a performance issue of CUDA kernel accessing GPU memory allocated by system malloc, more details are discussed in Section 4.4.3. If this performance issue can be resolved by NVIDIA, the performance of automatic offload can be further improved.
Large scale strong scaling tests were also performed for this code. Table 4 shows the runtime of CPU run, GPU run with native CUDA code, and automatic offload using Device First-Use policy. Scaling range goes from 25 nodes to 200 nodes. Throughout the tests, the automatic offload approach is consistently 2x faster than the native CUDA code, and up to 3x faster than the CPU run, breaking even with the extra node-hour charge rate for GPU, so users running on GPU not only get faster time-to-solution but also more efficiency in terms of energy consumption and cost. The strong scaling data is visualized in Figure 3, both the CPU run and automatic offload GPU run have excellent scability, reaching very close to linear scaling at this wide test range.